OpenStudy (anonymous):

Write the complex number in the form a + bi. 6(cos 330° + i sin 330°)

hero (hero):

Hint: Find the values of cos 330° and sin 330°

OpenStudy (anonymous):

\[\cos \left( 360-\theta \right)=\cos \theta\] \[\sin \left( 360-\theta \right)=-\sin \theta \]